# Burnside type results for fusion categories

### v10

@@ 11.1. Perfect Drinfeld center
We will characterize the perfect fusion categories with a perfect center. That will provide a large class of perfect modular fusion categories.
@@
Let $\mathcal{C}$ be a perfect fusion category. Then its universal grading group is perfect if and only if $\mathcal{Z}(\mathcal{C})$ is perfect.
@@
Let $\mathcal{C}$ be a braided fusion category. Then $\mathcal{Z}(\mathcal{C})$ is perfect if and only if $\mathcal{C}$ is perfect with a perfect universal grading group.
@@
Immediate by Proposition 11.2 because $\mathcal{C}$ is isomorphic to a fusion subcategory of $\mathcal{Z}(\mathcal{C})$. ∎
@@
Immediate by Theorem 7.8, because a grouplike element in the dual hypergroup of the Grothendieck ring of $\mathrm{Rep}(G)$ corresponds to a column of the character table of $G$ with squared norm $|G|$, so to a conjugacy class of size one, thus to a central element. ∎
@@
Let $G$ be a finite group. Then $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect if and only if $G$ is perfect with a trivial center $Z(G)$.
@@
Immediate by Corollary 11.3 and Lemma 11.4, because $\mathrm{Rep}(G)$ is perfect if and only if $G$ is perfect, and (as an abelian group) $Z(G)$ is perfect if and only if it is trivial. ∎
@@
Note that Corollary 11.5 can be proved alternatively using the fact (see [EGNO15, §8.5]) that the simple objects of $\mathcal{Z}(\mathrm{Rep}(G))$ are in bijection with the pairs $(C,V)$ where $C$ is a finite conjugacy class in $G$, and $V$ is an irreducible finite dimensional representation of the centralizer of $g$ in $C$, and the Frobenius-Perron dimension of the object corresponding to a pair $(C,V)$ is $|C|dim_{\mathbb{C}}(V)$.
@@
By Corollary 11.7, $\mathcal{Z}(\mathrm{Rep}(A_{5}))$ is a perfect integral modular fusion category of $\mathrm{FPdim}$ $360060^{2}=$ 3600. Using Remark 11.6 and GAP, it is of rank $22$ and type
$[[1,1],[3,2],[4,1],[5,1],[12,10],[15,4],[20,3]].$
@@
Is there a perfect integral modular fusion category of rank less than $22$? of $\mathrm{FPdim}$ less than $3600$?
Recall that a Hopf algebra $A$ is called perfect if $\mathrm{Rep}(A)$ is perfect, i.e. $G(A^{*})$ is a trivial group; that a finite dimensional semisimple Hopf algebra $A$ (over $\mathbb{C}$) is factorizable if and only if $\mathrm{Rep}(A)$ is modular; that $\mathcal{Z}(\mathrm{Rep}(G))=\mathrm{Rep}(D(G))$ where the Hopf algebra $D(G)$ is the Drinfeld double of $G$. So, for any non-abelian finite simple group $G$ then $D(G)$ is a finite dimensional semisimple factorizable perfect Hopf algebra (over $\mathbb{C}$) of dimension $|G|^{2}$.
@@
Is there a finite dimensional semisimple factorizable perfect Hopf algebra (over $\mathbb{C}$) of dimension less than $\dim(D(A_{5}))=3600$?
@@ 11.2. Burnside and integrality properties
Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring $K(\mathcal{C})$. By Theorem 7.8, the dual of $K(\mathcal{C})$ is perfect if and only if its universal grading group is trivial.
@@
The Grothendieck ring of a modular fusion category is perfect if and only if its dual is perfect (as it is isomorphic to its dual).
@@
By Remarks 11.11 and 11.12, a modular fusion category $\mathcal{C}$ is perfect if and only if its universal grading group is trivial (i.e. $\mathcal{C}_{ad}=\mathcal{C}$).
@@
We use the Galois action on $\widehat{\mathcal{B}}$ from Lemma 3.2. Let $\mu_{1}$ be the $\mathrm{FPdim}$ character. By Proposition 3.5, it follows that any Galois conjugate of $\mu_{1}$, say $\mu_{j}=\sigma.\mu_{1}$ has for formal codegree $n_{j}=\sigma(\mathrm{FPdim}(H,{\mathcal{B}}))$. On the other hand, since $\mu_{j}=\sigma.\mu_{1}$ is also a non-vanishing character, it should be a grouplike element, by the definition of a dual-Burnside hypergroup. Thus, its formal codegree should also be $\mathrm{FPdim}(H,{\mathcal{B}})$, which implies that $\sigma(\mathrm{FPdim}(H,{\mathcal{B}}))=\mathrm{FPdim}(H,{\mathcal{B}})$, for all $\sigma$ in the Galois group. Then, we are done since this implies that $\mathrm{FPdim}(H,{\mathcal{B}})$ is a rational number. ∎
The following result is partially due to Andrew Schopieray, see [Sch23].
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be rational abelian normalizable hypergroup. If the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect and Burnside then ${(H,\mathcal{B},\mu_{1})}$ is integral, i.e $d_{i}\in\mathbb{Q}$.
@@
The dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect, so as for the above proof of Theorem 1.10, $\sigma.\mu_{1}=\mu_{1}$ for all $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Thus $\sigma(d_{i})=d_{i}$, which shows that $d_{i}\in\mathbb{Q}$. ∎
@@
Recall that a modular fusion category is Burnside if and only if it is dual-Burnside (Corollary 10.3). Now, by Remarks 11.11 and 11.12, the universal grading group is trivial, so by Corollary 11.15, it must be integral. The converse follows from [GNN09, Theorem 6.1]. ∎
@@
By Theorem 1.23, a simple non-integral modular fusion category is not (dual-)Burnside. Plenty of modular fusion categories of Lie type (Verlinde categories) are simple non-integral, so is the one from [Sch22, Theorem 1], and the center of Extended-Haagerup, so none is (dual-)Burnside.
@@ 11.3. Near-group modular fusion categories
Let $G$ be a finite abelian group and $m$ be a non-negative integer. Let $K(G,m)$ be the fusion ring with basis $G\cup\{\rho\}$ and fusion rules:
@@
There is no modular fusion category whose Grothendieck ring is $K(G,m)$, with $G$ non-trivial and $m>0$.
@@
Let $\mathcal{C}$ be such a category and $R=K(\mathcal{C})$ be its Grothendieck ring. Since $\rho^{2}=m\rho+\sum_{g\in G}g$, then $\mathrm{FPdim}(\rho)^{2}=m\mathrm{FPdim}(\rho)+|G|$. Let $x_{1}>0$ and $x_{2}<0$ be the two solutions of the equation $x^{2}-mx-|G|=0$, then $\mathrm{FPdim}(\rho)=x_{1}$.
Any non-trivial character $\eta\in\widehat{G}$ can be extended to a character of $R$ (denoted also by $\eta$) by declaring $\eta(\rho)=0$. On the other hand, the trivial character $\eta=\epsilon$ ($\epsilon(g)=1$, for all $g\in G$) has two extensions on $R$: $\psi_{+}(\rho)=x_{1}$ and $\psi_{-}(\rho)=x_{2}$.
It follows that $\psi_{+}=\mathrm{FPdim}$. Thus
^=${\widehat{\mathcal{B}}}=\{\eta|\;\eta\in\widehat{G},\;\eta\neq\epsilon\}\cup\{\psi_{\pm}\}$
The multiplication on $\widehat{R}$ is given by
$\eta\star\psi_{-}=\eta,\;\eta\star\eta^{\prime}=\eta\eta^{\prime}\in\widehat{G},\;$
Moreover, it is easy to compute that:
$\psi_{-}^{2}=\frac{x_{2}^{2}+|G|}{x_{1}^{2}+|G|}\psi_{+}+\frac{x_{1}^{2}-x_{2}^{2}}{x_{1}^{2}+|G|}\psi_{-}.$
Thus $R$ has RN-dual since $x_{1}^{2}=mx_{1}+|G|>mx_{2}+|G|=x_{2}^{2}$.
Note that $R$ is a self dual hypergroup since $\mathcal{C}$ is modular. In particular, Drinfeld’s map $\mathrm{f}_{Q}:R\rightarrow\widehat{R}$, since an isomorphism of hypergroups, sends grouplike elements of $R$ to grouplike elements of $\widehat{R}$. Since $G(\widehat{R})=1$ and $G(R)=G$ is non-trivial, this shows that $\mathrm{f}_{Q}$ is not bijective, which is a contradiction. ∎
@@
Let $G$ be a finite group and $\mathcal{C}$ be a modular category whose Grothendieck ring ${K(\mathcal{C})}$ is $K(G,0)$. Then $G\simeq C_{2}$.
@@
As above, if $m=0$ then $\psi_{-}^{2}=\mathrm{FPdim}$ and therefore ${\widehat{{K(\mathcal{C})}}}$ has two grouplike elements $\psi_{+}=\mathrm{FPdim}$ and $\psi_{-}$. Note that Drinfeld’s map $\mathrm{f}_{Q}:{K(\mathcal{C})}\rightarrow{\widehat{{K(\mathcal{C})}}}$ send bijectively grouplike elements of ${K(\mathcal{C})}$ into grouplike elements of ${\widehat{{K(\mathcal{C})}}}$. Since all $g\in G$ are grouplike elements of ${K(\mathcal{C})}$ it follows $|G|=2$. ∎
@@
Ising’s category is the only example of a modular category whose Grothendieck ring is of type $K(C_{2},0)$, see [DGNO07, Appendix B].
@@
Immediate by Propositions 11.18, 11.19, Remark 11.20 and [Ost03b]. ∎
@@ 12. Concrete examples
This section focus on providing some concrete examples of fusion rings/categories to which the main results of the paper apply as categorification criteria, and also some counter-examples showing that some results cannot be generalized.
@@ 12.1. Burnside property
Recall that a fusion ring is called Burnside if for every basis element then its fusion matrix has norm $1$ if and only if its determinant is nonzero (in other words, grouplike is equivalent to invertible, see Proposition 4.3). Recall that William Burnside proved that the Grothendieck ring of $\mathrm{Rep}(G)$ is Burnside, for every finite group $G$, and that first author’s [Bur23, Theorem 2] generalizes this result to every weakly-integral fusion category (over $\mathbb{C}$).
A fusion ring $R$ is called $\alpha$-Frobenius if for all basis element $x$ then $\frac{\mathrm{FPdim}(R)^{\alpha}}{\mathrm{FPdim}(x)}$ is an algebraic integer. Kaplansky’s 6th conjecture states that the Grothendieck ring of a complex fusion category is $1$-Frobenius. Under the following bounds:
$\begin{array}[]{c|c|c|c|c|c|c|c|c}\text{rank}(R)&\leq 5&6&7&8&9&10&11&12\\ \hline\cr\mathrm{FPdim}(R)<&1000000&150000&15000&10000&4000&2000&500&200\end{array}$
there are exactly $192$ non-pointed simple integral $1$-Frobenius fusion rings with $\mathrm{FPdim}\neq p^{a}q^{b},pqr$ (the last restriction is motivated by [ENO11]). This is the current state of a classification initiated in [LPW21]. Among them, we found only $4$ non-Burnside ones. They have the type of $\mathrm{Rep}(A_{7})$, but different fusion data.
  • Rank: $9$,
@@
The determinant of the second fusion matrix (of each fusion data above) is ±$\pm 36$ whereas its norm is $6$, proving that these integral fusion ring are non-Burnside, and so cannot be categorified.
Similar classification were done in the non-1-Frobenius case, and unexpectedly, most (not all) of the ones we found are non-Burnside (see one example below), so without categorification, which can be seen as an evidence for Kaplansky 6th conjecture.
  • Rank: $7$,
  • FPdim: $798=2\times 3\times 7\times 19$,
  • Type: $[[1,1],[7,1],[8,1],[9,3],[21,1]]$,
@@
The determinant of the third fusion matrix is $16$ whereas its norm is $8$, proving that this integral fusion ring is non-Burnside, and so cannot be categorified.

### v11

@@ 11.1. Perfect Drinfeld center
We will characterize the perfect fusion categories with a perfect Drinfeld center. That will provide a large class of perfect modular fusion categories.
@@
Let $\mathcal{C}$ be a perfect fusion category. Then $\mathcal{Z}(\mathcal{C})$ is perfect if and only if the universal grading group of $\mathcal{C}$ is perfect.
@@
Let $\mathcal{C}$ be a braided fusion category. Then $\mathcal{Z}(\mathcal{C})$ is perfect if and only if $\mathcal{C}$ is perfect with a trivial universal grading group.
@@
If $\mathcal{Z}(\mathcal{C})$ is perfect, then $\mathcal{C}$ is also perfect because it can be embedded into $\mathcal{Z}(\mathcal{C})$ as a braided fusion category. According to Proposition 11.2, the universal grading group of $\mathcal{C}$ is perfect. However, since $\mathcal{C}$ is braided, its Grothendieck ring is commutative, which means its universal grading group must be abelian due to the faithful grading. But a perfect abelian group is trivial. Conversely, the result follows directly from Proposition 11.2, as the trivial group is perfect. ∎
Proposition 11.2 and Corollary 11.3 provide a complete characterization of when the Drinfeld center $\mathcal{Z}(\mathcal{C})$ is perfect assuming the fusion category $\mathcal{C}$ is perfect or braided. However, it is important to note that there are fusion categories that are neither perfect nor braided, yet their Drinfeld centers are perfect. An example of this is the fusion category $\mathcal{C}=\mathrm{Vec}(G)$, where $G$ is a non-abelian finite simple group. This fusion category is Morita equivalent to $\mathrm{Rep}(G)$, as illustrated in [EGNO15, Example 7.12.19]. As a result, their Drinfeld centers are braided equivalent, as shown in [EGNO15, Theorem 8.12.3]. Nonetheless, Corollary 11.7 below confirms that $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect in this case.
@@
Immediate by Theorem 7.8, because a grouplike element $\mu_{j}$ in the dual hypergroup of the Grothendieck ring of $\mathrm{Rep}(G)$ corresponds to a column of the character table of $G$ with squared norm $n_{j}=|G|$, by Lemma 2.31, so to a conjugacy class of size $|G|/n_{j}=1$, thus to a central element. ∎
@@
Let $G$ be a finite group. The Drinfeld center $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect if and only if $G$ is perfect with a trivial center $Z(G)$.
@@
Immediate by Corollary 11.3 and Lemma 11.4, because $\mathrm{Rep}(G)$ is perfect if and only if $G$ is perfect. ∎
@@
[…]
@@
By Corollary 11.7, $\mathcal{Z}(\mathrm{Rep}(A_{5}))$ is a perfect integral modular fusion category of $\mathrm{FPdim}$ $60^{2}=3600$. Using Remark 11.6 and GAP, it is of rank 222222 and type
[[1,1],[3,2],[4,1],[5,1],[12,10],[15,4],[20,3]].[[1,1],[3,2],[4,1],[5,1],[12,10],[15,4],[20,3]].[ [ 1 , 1 ] , [ 3 , 2 ] , [ 4 , 1 ] , [ 5 , 1 ] , [ 12 , 10 ] , [ 15 , 4 ] , [ 20 , 3 ] ] .
@@
Is there a non-trivial perfect integral modular fusion category of rank less than 222222, or $\mathrm{FPdim}$ of less than 360036003600?
Recall that a Hopf algebra $A$ is called perfect if $\mathrm{Rep}(A)$ is perfect, i.e. $G(A^{*})$ is a trivial group; that a finite dimensional semisimple Hopf algebra $A$ (over $\mathbb{C}$) is factorizable if and only if $\mathrm{Rep}(A)$ is modular; that $\mathcal{Z}(\mathrm{Rep}(A))=\mathrm{Rep}(D(A))$ where the Hopf algebra $D(A)$ is the Drinfeld double of $A$. So, for any non-abelian finite simple group $G$ then $D(G)$ is a finite dimensional semisimple factorizable perfect Hopf algebra (over $\mathbb{C}$) of dimension $|G|^{2}$. Let us clarify Question 11.8 for the case of Hopf algebras:
@@
Is there a non-trivial finite-dimensional semisimple factorizable perfect Hopf algebra over $\mathbb{C}$ that has less than 222222 irreducible representations (up to isomorphism), or a dimension of less than 360036003600?
@@ 11.2. Burnside and integrality properties
According to Theorem 7.8, the dual hypergroup of a commutative fusion ring is perfect if and only if the universal grading group of the fusion ring is trivial.
@@
The Grothendieck ring of a modular fusion category is perfect if and only if its dual is perfect (since it is isomorphic to its dual as normalized hypergroup, see §10.1).
@@
By Remarks 11.11 and 11.12, a modular fusion category $\mathcal{C}$ is perfect if and only if its universal grading group is trivial (i.e. $\mathcal{C}_{ad}=\mathcal{C}$). Alternatively, recall that $({\mathcal{C}_{\mathrm{pt}}})^{\prime}={\mathcal{C}_{\mathrm{ad}}}$ by [EGNO15, Corollary 8.22.8].
@@
We utilize the Galois action on $\widehat{\mathcal{B}}$ as described in Lemma 3.2. Let $\mu_{1}$ be the $\mathrm{FPdim}$ character. According to Proposition 3.5, any Galois conjugate of $\mu_{1}$, denoted by $\mu_{j}=\sigma\cdot\mu_{1}$, has a formal codegree given by $n_{j}=\sigma(\mathrm{FPdim}(H,\mathcal{B}))$. However, because $\mu_{j}=\sigma\cdot\mu_{1}$ is also a non-vanishing character, it must be a grouplike element due to the definition of a dual-Burnside hypergroup (Definition 4.7). Therefore, its formal codegree must also be $\mathrm{FPdim}(H,\mathcal{B})$, by Corollary 2.29. Thus, $\sigma(\mathrm{FPdim}(H,\mathcal{B}))=\mathrm{FPdim}(H,\mathcal{B})$ for every $\sigma$ in the Galois group. It follows that $\mathrm{FPdim}(H,\mathcal{B})$ is a rational number, completing the proof. ∎
As a consequence of Theorem 1.13, every commutative dual-Burnside fusion ring is weakly integral. However, the converse is not true, even for unitary integral fusion categories such as $\mathrm{Rep}(A_{7})$; refer to §12.2 for more examples. The following result is partially attributed to Andrew Schopieray; see [Sch23].
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be rational abelian normalizable hypergroup. If the dual $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect and Burnside then ${(H,\mathcal{B},\mu_{1})}$ is integral, i.e $d_{i}\in\mathbb{Q}$.
@@
The dual $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect, meaning that $\mu_{1}$ is the only grouplike linear character. As in the proof of Theorem 1.13, for any $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, $\sigma.\mu_{1}$ is also non-vanishing and thus grouplike by the dual-Burnside assumption. Consequently, $\sigma.\mu_{1}=\mu_{1}$, and $\sigma(d_{i})=d_{i}$ for all $\sigma$, which implies that $d_{i}\in\mathbb{Q}$. ∎
@@
Recall that a modular fusion category is Burnside if and only if it is dual-Burnside (Corollary 10.3). Now, by Remark 11.13, the universal grading group is trivial, so by Corollary 11.16, it must be integral. The converse follows from [GNN09, Theorem 6.1]. ∎
@@
According to Theorem 1.27, a simple non-integral modular fusion category cannot be (dual-)Burnside. Many modular fusion categories of Lie type, known as Verlinde categories, belong to this class, as do the one described in [Sch22, Theorem 1] and the Drinfeld center of the Extended-Haagerup fusion categories. Consequently, none of these are (dual-)Burnside. Therefore, for the Grothendieck rings in this class, we can still infer the existence of a non-group-like and non-vanishing linear character (and basic element).
@@ 11.3. Near-group modular fusion categories
Let $G$ be a finite abelian group and $m$ be a non-negative integer. Let $K(G,m)$ be the fusion ring with basis $\mathcal{B}=G\cup\{\rho\}$ and fusion rules:
@@
There is no modular fusion category of Grothendieck ring $K(G,m)$ if $G$ is non-trivial and $m>0$.
@@
Let $\mathcal{C}$ be a modular fusion category of Grothendieck ring $K(G,m)$. Given that $\rho^{2}=m\rho+\sum_{g\in G}g$, it follows that $\mathrm{FPdim}(\rho)^{2}=m\mathrm{FPdim}(\rho)+|G|$. Let $x_{+}>0$ and $x_{-}<0$ be the solutions to the equation $x^{2}-mx-|G|=0$. Thus, $\mathrm{FPdim}(\rho)=x_{+}$. Now, consider $\eta$, a linear character of $R$. For all $g\in G$, we have $\eta(\rho)=\eta(\rho g)=\eta(\rho)\eta(g)$. Therefore, $\eta(\rho)$ is nonzero if and only if $\eta(g)=1$ for all $g\in G$. This condition is equivalent to solving $\eta(\rho)^{2}-m\eta(\rho)-|G|=0$, which implies $\eta(\rho)=x_{\pm}$. We denote such characters as $\psi_{\pm}$. It follows that $\psi_{+}=\mathrm{FPdim}$ is the unit of $^{\widehat{\mathcal{B}}}$. Let $\epsilon$ be the trivial character of $G$. We identify any $\eta$ in $\widehat{G}\setminus\{\epsilon\}$ (non-empty since $G$ is non-trivial) with the corresponding element in $^{\widehat{\mathcal{B}}}$ that vanishes on $\rho$. Let $q=-\frac{x_{-}}{x_{+}}$, which is positive. The multiplication on $\widehat{R}$ is defined by:
$\eta\star\psi_{-}$ $=\eta,$
$\eta\star\eta^{\prime}$ $=(1-\delta_{\eta^{\prime},\eta^{*}})\eta\eta^{\prime}+\delta_{\eta^{\prime},\eta^{*}}(\frac{q}{1+q}\psi_{+}+\frac{1}{1+q}\psi_{-}),$
$\psi_{-}\star\psi_{-}$ $=q\psi_{+}+(1-q)\psi_{-}.$
[…]
@@
Let $\mathcal{C}$ be a modular fusion category of Grothendieck ring $K(G,0)$. Then $G\simeq C_{1}$ or $C_{2}$.
@@
[…]
@@
Ising categories are the only modular fusion categories whose Grothendieck ring are $K(C_{2},0)$, see [DGNO10, Appendix B].
@@
According to Proposition 11.18, either $G$ is trivial or $m=0$. If $G$ is trivial, the category has rank 2, and the conclusion follows from [Ost03b]. On the other hand, if $G$ is non-trivial and $m=0$, the result follows from Proposition 11.19 and Remark 11.20. ∎
@@ 12. Concrete examples
This section focuses on providing concrete examples of fusion rings and categories to which the main results of the paper apply as categorification criteria, along with counter-examples demonstrating the limits of these results.
@@ 12.1. Burnside property
Recall that a fusion ring is called Burnside if, for every basis element, its fusion matrix has a norm of 111 if and only if its determinant is nonzero. In other words, being grouplike is equivalent to being invertible (see Proposition 4.3). William Burnside originally proved that the Grothendieck ring of $\mathrm{Rep}(G)$ is Burnside for every finite group $G$. This result was expanded in [Bur23] to include all weakly integral fusion categories with a commutative Grothendieck ring, but first within the context of modular categories in [GNN09, Appendix]. In Theorem 1.2, it extends to every commutative fusion ring with an $h$-integral dual, and further to hypergroups in Theorem 4.17.
Remark 12.1.
By Theorems 1.8 and 7.9, along with the fact that $\mathrm{Rep}(G)_{pt}\simeq\mathrm{Rep}(G/G^{\prime})$, Burnside’s result can be restated as follows: for any finite group $G$,
$\left(\prod_{j\in\mathcal{I}}\frac{C_{j}}{|\mathcal{C}^{j}|}\right)^{2}=\frac{1}{|G^{\prime}|}\sum_{\mathcal{C}^{j}\subseteq G^{\prime}}C_{j},$
where $G^{\prime}$ is the commutator subgroup of $G$, $(\mathcal{C}^{j})$ are the conjugacy classes of $G$, and $C_{j}:=\sum_{g\in\mathcal{C}^{j}}g$ is the class sum associated with $\mathcal{C}^{j}$. It was already observed in [Har08].
A fusion ring $R$ is termed $\alpha$-Frobenius if, for every basis element $x$, the expression $\frac{\mathrm{FPdim}(R)^{\alpha}}{\mathrm{FPdim}(x)}$ is an algebraic integer [ENO25]. Kaplansky’s 6th conjecture posits that the Grothendieck ring of a complex fusion category is 111-Frobenius [ENO11, Question 1]. An exhaustive classification of all 111-Frobenius simple integral fusion rings, within certain specified limits, was provided in [LPW21]. These limits have been updated in [BP24] as follows:
$\begin{array}[]{c|cccccccc}\text{Rank}&\leq 5&6&7&8&9&10&11&12\\ \hline\cr\operatorname{\mathsf{FPdim}}\leq&10^{7}&10^{6}&10^{5}&20000&10000&5000&3000&1000\end{array}$
There are exactly 505505505 non-pointed examples (including 888 that are character rings of groups). Among them, we found only 444 non-Burnside ones. They have the type of $\mathrm{Rep}(A_{7})$, but different fusion data.
  • Rank: 999,
@@
The determinant of the second fusion matrix for each fusion data set mentioned above is ±36$\pm$ 36± 36, while its norm is 666. This indicates that these integral fusion rings are non-Burnside and, consequently, cannot be categorified.
Similar classifications have been conducted for the non-1-Frobenius case. Interestingly, most—but not all—of the fusion rings identified are non-Burnside, and thus, not amenable to categorification (see the example below). This serves as supporting evidence for Kaplansky’s 6th conjecture.
  • Rank: 777,
  • FPdim: 798=2×3×7×19798=2$\times$ 3$\times$ 7$\times$ 19798 = 2 × 3 × 7 × 19,
  • Type: [[1,1],[7,1],[8,1],[9,3],[21,1]][[1,1],[7,1],[8,1],[9,3],[21,1]][ [ 1 , 1 ] , [ 7 , 1 ] , [ 8 , 1 ] , [ 9 , 3 ] , [ 21 , 1 ] ],
@@
The determinant of the third fusion matrix is 161616, while its norm is 888. This shows that the integral fusion ring is non-Burnside and therefore cannot be categorified.